OpenStudy (gabylovesyou):

Find the area of a circle with a diameter of 16 inches. Use 3.14 for pi. 50.24 in2 100.48 in2 200.96 in2 251.2 in2

jimthompson5910 (jim_thompson5910):

Use A = pi*r^2

jimthompson5910 (jim_thompson5910):

keep in mind that radius = (diameter)/2

OpenStudy (gabylovesyou):

i get 201.06

jimthompson5910 (jim_thompson5910):

how are you getting that?

OpenStudy (gabylovesyou):

diameter is 16.... 16 / 2 = 8 8 * 8 = 64 64 * 3.14 = 200.96 (nevermind) lol

OpenStudy (goformit100):

We know Area of circle is  Solve it, how much you got ?